### v1

# Lightweight Prompt Biasing for Contextualized End-to-End ASR Systems

Abstract
End-to-End Automatic Speech Recognition (ASR) has advanced significantly yet still struggles with rare and domain-specific entities. This paper introduces a simple yet efficient prompt-based biasing technique for contextualized ASR, enhancing recognition accuracy by leverage a unified multi-task learning framework. The approach comprises two key components: a prompt biasing model which is trained to determine when to focus on entities in prompt, and a entity filtering mechanism which efficiently filters out irrelevant entities. Our method significantly enhances ASR accuracy on entities, achieving a relative 30.7% and 18.0% reduction in Entity Word Error Rate compared to the baseline model with shallow fusion on in-house domain dataset with small and large entity lists, respectively. The primary advantage of this method lies in its efficiency and simplicity without any structure change, making it lightweight and highly efficient.
Index Terms:
speech recognition, contextual biasing, human-computer interaction
I Introduction
In recent years, End-to-End (E2E) automatic speech recognition (ASR) systems have made significant progress thanks to advances in deep learning models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, these systems still face challenges in accurately recognizing rare words and domain-specific terms. To enhance ASR performance, researchers have proposed various methods [11, 12, 13, 14, 15] to improve the performance on specific contexts. Contextual biasing techniques leverage external information to boost recognition accuracy, particularly when handling rare or specialized entities. Typically, the external information is provided to the ASR system during decoding in the form of an words/entities list, which is often referred as biasing list.
Extensive research has been conducted on contextual biasing techniques for E2E ASR systems, which can be broadly classified into two major categories. Shallow fusion methods [11, 12] integrate external language models (LMs) with ASR systems during decoding by weighting LM scores, allowing the system to prioritize contextually relevant terms. Further advancements in shallow fusion include sub-word regularization, pre-training, grapheme-to-grapheme pronunciation learning, and deep integration with neural network language models [16, 17]. However, a fundamental limitation of shallow fusion is its reliance on post-contextual boosting, which requires the model to generate the correct expected prefix terms in the candidate pool without access to external contextual information.
To address these limitations, deep biasing methods directly incorporate contextual information into E2E ASR models, enabling joint optimization. A notable example is the Contextual Listen, Attend and Spell (CLAS) system, which integrates ASR components with contextual embeddings [13]. The CLAS system has been further enhanced to include phonetic information, leveraging pronunciation knowledge for improved recognition of rare words [18, 19]. Other approaches have explored intermediate biasing loss and attention mechanisms to improve contextual modeling in CLAS systems [20]. However, these methods often rely on additional encoders to embed contextual information, introducing computational overhead and complexity. Moreover, attention mechanisms may struggle to scale effectively with large biasing lists, a common scenario in real-world applications [13].
Transformer-based architectures have recently become the backbone of state-of-the-art E2E ASR systems [21]. These models consistently outperform traditional Recurrent Neural Network (RNN) approaches and are now widely used in both research and industry [1, 22, 23, 24, 25, 2, 26]. However, integrating contextual biasing effectively into Transformer models remains a challenging problem, with direct impact on user experience and commercial deployment.
There is also increasing interest in utilizing large language models (LLMs) to enhance contextual biasing in ASR [27, 28]. Although these methods have demonstrated promising results, they often require significant computational resources and add architectural complexity. As LLM-based solutions are still emerging and not yet widely adopted in production ASR systems, our work focuses on practical Transformer-based E2E ASR models, leaving LLM integration for future research.
In this work, we propose a straightforward and effective method for contextual biasing in Transformer-based ASR systems using a multi-task learning framework. Our approach, called Prompt Biasing, leverages the Transformer’s cross-attention mechanism to introduce contextual information as a prompt to the decoder. By employing a unified multi-task learning setup with dedicated task tokens, the model can efficiently handle both biasing and non-biasing scenarios without requiring any architectural changes. Additionally, we introduce an efficient entity filtering strategy that rely on the same model during decoding, enabling robust performance even with large biasing lists. Experimental results demonstrate that our lightweight Prompt Biasing model consistently outperforms the baseline with shallow fusion on domain-specific datasets, achieving a relative reduction in Entity Word Error Rate (EWER) of 30.7% and 18.0% for small and large biasing lists, respectively.
(a) Standard Transformer ASR training
(b) Unified multi-task training for contextual biasing
(c) Multi-task token format for Prompt Biasing training
Figure 1: Overview of the proposed Prompt Biasing method. (a) Standard Transformer ASR training, where the model predicts the next token from previous tokens. (b) Unified multi-task training for contextual biasing, where the model predicts the next token using both previous tokens and contextual information provided as a prompt. (c) Multi-task token format for Prompt Biasing training: biasing and non-biasing tasks are indicated by special task tokens (<hit>/<miss>), and the biasing list is included as a prompt starting with the <sop> token.
II Related Work
This section provides an overview of recent advancements in Transformer-based ASR systems and contextual biasing techniques, with a particular focus on state-of-the-art developments in Transformer architectures.
II-A Transformers based E2E ASR
Recent advancements in Transformer-based architectures have significantly advanced the field of end-to-end ASR. The Speech-Transformer [23] first demonstrated the effectiveness of self-attention and cross-attention mechanisms for modeling long-range dependencies in speech, resulting in improved recognition accuracy and efficiency. Building on this, the Transformer-Transducer framework incorporates a Transformer encoder within a transducer model, enabling efficient streaming recognition with competitive performance [26, 29]. The Conformer architecture further enhances Transformer models by integrating convolutional layers, which capture local acoustic features and improve robustness in challenging acoustic conditions [24]. Collectively, these innovations have established Transformer-based models as the state-of-the-art for ASR, offering superior accuracy, scalability, and computational efficiency compared to traditional RNN-based systems.
More recently, Whisper [25] has emerged as a versatile end-to-end Transformer-based speech processing system, trained on large-scale multilingual data. Whisper supports not only speech recognition but also translation and language identification within a unified framework, further demonstrating the flexibility and effectiveness of Transformer architectures in speech applications.
II-B Contextual biasing
While shallow fusion techniques can be easily applied to Transformer-based ASR systems, their performance improvements remain limited due their post-training adjustment, and also be concerned to either under- or over- biasing issue [11]. Deep biasing techniques, such as the CLAS system [13], have historically relied on RNN architectures.
Some studies have attempted to enhance contextual biasing capabilities in Transformer-based models. For instance, certain works introduce a Tree-Constrained Pointer Generator (TCPGen) module [30], which dynamically adjusts transcription by interpolating between the original model and TCPGen distributions [31]. Another approach, PromptASR, integrates an additional pre-trained text encoder and injects contextual information by adding cross-attention modules after the acoustic self-attention modules[32]. Authors in [33] proposed to introduce contextual adapters for personalization in neural transducer based ASR models.These methods either introduce auxiliary models or require significant architectural modifications to improve contextual biasing in Transformer ASR models. In contrast, our approach is designed to avoid any architectural modifications. We leverage the native multi-task learning capabilities of the Transformer model, enabling effective contextual biasing without introducing extra computational overhead or complexity.
III Proposed Approach
In this section, we outline the core principles of integrating biasing tasks into ASR systems via a multi-task learning framework, and we further detail the entity filtering technique employed during the decoding process.
III-A Multi-task Learning for Contextual Biasing
Figure 1a illustrates a standard Transformer-based ASR system focused solely on speech recognition. In contrast, Whisper [25] introduces a unified multi-task framework using conditional prefix tokens (e.g., transcription or translation) to differentiate tasks during training and decoding. However, Whisper does not natively support contextual biasing functionality.
Inspired by Whisper [25], we propose to formulate the contextual biasing task as a multi-task learning problem, where the model is trained to handle both biasing and non-biasing tasks within a unified framework. The proposed multi-task learning framework is illustrated in Figure 1b. In this framework, the model is trained to predict the next token based on both the previous tokens and the contextual information provided as a prompt. The contextual information is typically a list of entities or phrases that the model should focus on during recognition. This prompt is integrated into the Transformer decoder, allowing the model to condition its predictions on the provided contextual information.
As illustrated in Figure 1c, we adopt a unified training format for both biasing and non-biasing tasks. Each training sample is structured similarly, with special tokens indicating the presence or absence of entities. The <hit> token marks entities that appear in the audio, while the <miss> token marks those that do not. To clearly separate the prompt (i.e. contextual information) from the transcription, we use a special <sop> token at the start of the prompt and the standard <sot> token at the beginning of the transcription. When no contextual information is provided (i.e., for regular recognition), the <miss> token is used, treating it as a non-biasing task.
The core innovation of our approach lies in employing multi-task learning with specialized task tokens to distinguish between biasing and non-biasing tasks below:
  1. 1.
    Biasing Task: When certain entities in the prompt appears in the audio, the model is given the <hit> token. This task token directs the model to concentrate on the prompt content, thereby effectively leveraging the external contextual information to improve the recognition accuracy on these hit entities. Not all entities in the prompt are required to be present in the audio, and the model is trained to learn where to focus from data.
  2. 2.
    Non-biasing Task: When none of entities in the prompt are present in the audio, the model receives the <miss> token. This task token instructs the model to disregard the prompt, enhancing its robustness by reducing the influence of irrelevant information.
The introduction of specialized task tokens enables our multi-task framework to seamlessly handle both biasing and non-biasing tasks within a unified architecture. During inference, these tokens allow the model to accurately distinguish when contextual prompt information should be utilized and when it should be disregarded, ensuring robust and context-appropriate recognition performance.
TABLE I: Examples from models trained With/Without task tokens. The biasing list includes coffee, milk, chocolate.
System Task Tokens Output
Reference N/A So I’ve added four toffee almond milk hot cocoas what else?
Recognition Without coffee coffee milk hot cocoa no no no no no
Recognition With So I’ve added 4 toffee almond milk hot cocos what else?
Table I presents outputs from models trained with and without task tokens. The absence of task tokens leads the model to overfit to the prompt, resulting in repeated or irrelevant words (e.g., coffee) and reduced transcription accuracy. In contrast, incorporating task tokens enables the model to correctly identify and utilize relevant entities from the prompt while preserving accurate recognition of the remaining audio content. This example highlights that merely supplying prompt information is insufficient; without a multi-task framework and explicit task tokens, the model is prone to hallucinations and transcription errors. The proposed multi-task approach with task tokens allows the model to effectively differentiate between biasing and non-biasing scenarios, leveraging contextual information only when appropriate. This design enhances transcription accuracy and facilitates the integration of contextual biasing without introducing architectural complexity.
Algorithm 1 Entity Filtering
0:  Audio input, Entities in biasing list
0:  Filtered list of entities
1:  $hi$: index of the <hit> token
2:  $H\leftarrow\text{Encoder}(\text{audio})$Compute encoder embedding}
3:  for each entity $E$ in the biasing list do
4:     $L\leftarrow\text{Decoder}(E,H)$Compute logits for task tokens}
5:     $P_{h}\leftarrow\text{softmax}(L)[hi]$Probability on <hit> token}
6:     $P_{\text{hit}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}P_{h}[i]$Average over sub-word units}
7:  end for
8:  Filter out entities with $P_{\text{hit}}<0.5$
9:  return  Filtered list of entities
Figure 2: Decoding process for the Prompt Biasing model. The model is trained to predict <hit> or <miss> for each sub-word in the prompt, enabling effective filtering of irrelevant entities. During decoding, entity filtering (see Algorithm 1) greatly reduces the biasing list size.
III-B Entity Filtering for Biasing
Previous studies on CLAS have shown that attention-based neural biasing methods encounter difficulties when processing large biasing lists [13]. Our own experiments, as reflected in rows C1–C3 of Table II, corroborate this finding: large biasing lists introduce substantial noise, which diminishes the model’s ability to accurately focus on relevant entities. Our analysis reveals a clear correlation between the precision of the biasing list and the effectiveness of contextual biasing—the more targeted and relevant the biasing list, the greater the improvement in biasing performance. This observation motivates the introduction of an entity filtering mechanism to systematically eliminate irrelevant candidates from large biasing lists.
In our approach, we utilize the multi-task trained model to efficiently filter out irrelevant entities from large biasing lists. Unlike typical Transformer training schemes, where losses on conditional input tokens are ignored and these tokens serve solely as contextual cues, our method treats prompt tokens as both conditional and predictive inputs. Specifically, during training, each sub-word unit of the entities in the prompt is labeled with either a <hit> or <miss> token, depending on its presence in the audio content (see Figure 1b). This enables the model to learn to identify which entities in the prompt are actually present in the audio. Consequently, the same model can be directly leveraged for entity filtering during inference, eliminating the need for a separate filtering model.
As depicted in Figure 2, we adopt a progressive entity filtering strategy during decoding, utilizing the multi-task model to identify and exclude irrelevant entities from the biasing list. This process substantially reduces the size of the candidate list; for example, initial lists containing up to 2,000 entities can be efficiently narrowed to 10–20 highly relevant candidates. By eliminating distractors, the filtering mechanism enhances the model’s ability to focus on pertinent entities, thereby improving recognition accuracy. The detailed steps of this entity filtering approach are provided in Algorithm 1. Although the method requires evaluating each entity in the biasing list, it is highly amenable to parallelization and can be executed with minimal latency on modern hardware accelerators such as GPUs.
TABLE II: Performance on in-house domain dataset with various biasing lists.
Row Model Biasing List EWER(%) WER(%)
A1 Baseline N/A 7.76 4.75
B1 Baseline + Shallow Fusion Small 5.83 4.67
B2 Baseline + Shallow Fusion Large 6.04 4.74
C1 Prompt Biasing Exact 1.80 4.60
C2 Prompt Biasing Small 4.16 4.54
C3 Prompt Biasing Large 5.61 4.80
D1 Prompt Biasing + Entity Filtering Small 4.04 4.46
D2 Prompt Biasing + Entity Filtering Large 4.95 4.71
IV Experiments
Our model utilizes a Transformer encoder-decoder architecture, with the encoder enhanced by Conformer layers [24] to improve speech recognition accuracy. To increase computational efficiency, a convolutional down-sampling module is applied prior to the Conformer layers, reducing the input frame rate by a factor of 8. The encoder consists of 18 layers, while the decoder comprises 6 layers. The model is initially pre-trained on a comprehensive in-house dataset encompassing diverse audio conditions and interaction scenarios, following the standard speech recognition configuration shown in Figure 1a. This pre-trained model serves as the baseline for all subsequent Prompt Biasing experiments, enabling a direct assessment of the proposed biasing approach.
IV-A Training Details
We fine-tune our model from a pre-trained Transformer backbone, eliminating the need for training from scratch. The vocabulary is augmented with special tokens (<sop>, <hit>, and <miss>) to support the proposed multi-task learning framework. The training dataset comprises both biasing and non-biasing samples, with biasing samples constituting 65% of the total. For each sample, entities in the biasing list are randomly selected from either the reference transcription or a pool of negative phrases drawn from diverse text corpora, with each entity limited to a maximum of five words. Task tokens are assigned based on the presence of these entities in the audio. To manage memory consumption, each training sample contains fewer than 20 entities in the biasing list. The model is trained on approximately 5,400 hours of anonymized paired audio-text data using the AdamW optimizer [34] and a linear decay learning rate schedule, with a peak learning rate of $2.24\times 10^{-4}$ and a brief warmup phase.
IV-B Evaluation Settings
We conducted a comprehensive evaluation of the proposed Prompt Biasing approach using an in-house domain-specific dataset covering 9 popular domains and comprising approximately 570,000 words. Each utterance was annotated by human experts to identify entities of interest. To measure recognition performance on these entities, we utilized the Entity Word Error Rate (EWER), which specifically quantifies errors within labeled entities.
To systematically investigate the effect of biasing list size, we constructed three types of biasing lists for each utterance:
  1. 1.
    Exact: Consists solely of the entities present in the reference transcription for the given utterance.
  2. 2.
    Small: Augments the Exact list with approximately 50 randomly selected distractor entities.
  3. 3.
    Large: Augments the Exact list with approximately 1800 randomly selected distractor entities.
To assess the robustness of the model to noisy contextual information, we conducted experiments on a large in-house dataset comprising 7.6 million words. For each utterance, approximately 100 entities were randomly sampled from external text sources and added as distractors to the prompt, simulating the presence of irrelevant entities. The overall Word Error Rate (WER) was then evaluated on this dataset to determine the model’s resilience to such noise.
For a fair comparison with conventional contextual biasing methods in E2E ASR systems, we employ the shallow fusion technique with the baseline model. During decoding, the provided biasing list is dynamically converted into a weighted finite-state transducer (WFST) graph, following the approach described in [11]. This WFST is then integrated into the baseline model’s beam search process. The shallow fusion weight is set to 1.6 to achieve a balance between contextual biasing effectiveness and overall transcription accuracy.
IV-C Experimental Results
IV-C1 Contextual Biasing Performance
Rows C1–C3 in Table II present the performance of the Prompt Biasing model evaluated with three different types of biasing lists on our in-house domain dataset. First of all, when an ideal ground truth list (i.e. Exact list) is provided, the Prompt Biasing model achieves an impressive EWER of 1.8%, significantly outperforming its baseline model (i.e. A1 in Table II) by 76.8% relatively. This demonstrates the effectiveness of our approach in leveraging contextual information to enhance the entity recognition accuracy. In addition to the strong EWER performance, the Prompt Biasing model achieves a competitive WER of 4.60%, slightly outperforming the baseline. This indicated that the model effectively leverages contextual information to improve entity recognition while maintaining overall transcription quality. In addition to evaluating the Prompt Biasing model with the ideal Exact biasing list, we also assess its performance using two more realistic, noisy biasing lists: Small and Large. These lists better reflect practical scenarios where the biasing list may not be entirely accurate or precise. The Small list contains approximately 50 entities, while the Large list includes around 1800 entities. As shown in Table II (rows C2 and C3), the Prompt Biasing model achieves strong EWERs of 4.16% and 5.61% with the Small and Large lists, respectively, along with competitive WERs of 4.54% and 4.80%. These results demonstrate that the model maintains effective contextual biasing and robust general transcription accuracy even when provided with imperfect biasing lists.
Beyond the strong EWER performance of the Prompt Biasing model across various biasing lists, we observe a clear trend of increasing EWER as the size of the biasing list grows. Specifically, EWER rises from 1.80% with the Exact list to 4.16% with the Small list, and further to 5.61% with the Large list. This pattern suggests that larger and noisier biasing lists introduce additional challenges, reducing the model’s ability to accurately focus on relevant entities. These results underscore the importance of providing precise and relevant contextual information to maximize the effectiveness of contextual biasing.
IV-C2 Impact of Entity Filtering
Given the observed benefits of providing a precise biasing list, we further evaluate the impact of our entity filtering strategy detailed in Algorithm 1. As shown in rows D1 and D2 of Table II, applying entity filtering leads to a notable improvement in EWER for both the Small and Large noisy biasing lists, compared to the standard prompt biasing model (rows C2 and C3). Specifically, EWER decreases from 4.16% to 4.04% for the Small list and from 5.61% to 4.95% for the Large list. These results demonstrate that entity filtering effectively enhances biasing performance, particularly when handling large and noisy biasing lists. Additionally, general transcription performance, as measured by WER, is also slightly improved—from 4.54% to 4.46% for the Small list and from 4.80% to 4.71% for the Large list—indicating that the entity filtering strategy not only improves entity recognition but also helps maintain overall transcription quality.
IV-C3 Comparison to Shallow Fusion
Shallow fusion is a widely adopted technique for enhancing contextual biasing performance in ASR systems. As shown in Table II, we compare our proposed Prompt Biasing model with entity filtering (rows D1 and D2) against the baseline model with shallow fusion (rows B1 and B2) across different biasing list sizes. Our approach consistently outperforms shallow fusion, achieving a relative reduction in EWER of 30.7% for the Small biasing list (D1 vs. B1) and 18.0% for the Large biasing list (D2 vs. B2). These results highlight the effectiveness of Prompt Biasing with entity filtering in leveraging contextual information to improve entity recognition accuracy. Additionally, our method maintains competitive overall transcription quality, with WERs of 4.46% and 4.71% for the Small and Large lists, respectively, compared to 4.67% and 4.74% for shallow fusion. This demonstrates that our approach not only enhances entity recognition but also preserves general ASR performance, even with large and noisy biasing lists.
IV-C4 Robustness Against Noise
Table III presents the performance of the models on a 7.6M-word in-house dataset under both standard (no biasing list) and noisy biasing list conditions. The standard setting corresponds to conventional ASR without contextual information. Compared to the baseline, the Prompt Biasing model exhibits only a marginal increase in WER of 0.04% in the absence of contextual information, demonstrating minimal impact on general ASR performance. Furthermore, when evaluated with a noisy biasing list, the Prompt Biasing model shows a negligible WER degradation of 0.06%, indicating strong robustness to irrelevant or inaccurate contextual information.
TABLE III: Performance on in-house dataset for robustness benchmark.
Model Biasing List WER(%)
Baseline N/A 6.91
Prompt Biasing N/A 6.95
Prompt Biasing Noisy 6.97
V Conclusion and Future Work
In this paper, we presented a lightweight and effective prompt-based contextual biasing approach for Transformer-based ASR systems. By casting contextual biasing as a multi-task learning problem, our method enables seamless integration of contextual information without requiring any architectural modifications. We also introduced an efficient entity filtering mechanism that significantly enhances biasing performance, especially when dealing with large and noisy biasing lists. Experimental results show that our approach consistently outperforms shallow fusion techniques in both entity recognition and overall transcription accuracy. In future work, we plan to investigate the combination of Prompt Biasing with shallow fusion and explore the integration of large audio language models to further advance contextual ASR capabilities.

### v2

@@
 During inference, these tokens allow the model to accurately distinguish when contextual prompt information should be utilized and when it should be disregarded, ensuring robust and context-appropriate recognition performance.
-TABLE I: Examples from models trained With/Without task tokens. The biasing list includes coffee, milk, chocolate.
-System Task Tokens Output
-Reference N/A So I’ve added four toffee almond milk hot cocoas what else?
-Recognition Without coffee coffee milk hot cocoa no no no no no
-Recognition With So I’ve added 4 toffee almond milk hot cocos what else?
-Table I presents outputs from models trained with and without task tokens.
-The absence of task tokens leads the model to overfit to the prompt, resulting in repeated or irrelevant words (e.g., coffee) and reduced transcription accuracy.
-In contrast, incorporating task tokens enables the model to correctly identify and utilize relevant entities from the prompt while preserving accurate recognition of the remaining audio content.
+TABLE I: WERs of models trained With/Without task tokens on a food ordering dataset with/without biasing list.
+Model trained
+with Task Tokens
+w/o biasing list w/ biasing list
+No 10.11 19.21
+Yes 9.70 9.46
+TABLE II: Examples from models trained With/Without task tokens. The biasing list includes coffee, milk, chocolate.
+System Result
+Reference
+So I’ve added four toffee almond milk
+hot cocoas what else?
+Model Without
+Task Tokens
+coffee coffee milk
+hot cocoa no no no no no
+Model With
+Task Tokens
+So I’ve added 4 toffee almond milk
+hot cocos what else?
+Figure 2: Decoding process for the Prompt Biasing model.
+The model is trained to predict <hit> or <miss> for each sub-word in the prompt, enabling effective filtering of irrelevant entities.
+During decoding, entity filtering (see Algorithm 1) greatly reduces the biasing list size.
+Table I summarizes initial results from an early investigation using a food ordering dataset, comparing models trained with and without the inclusion of specialized task tokens.
+The models were assessed under two settings: with a biasing list and without one (i.e., standard ASR).
+The findings reveal that models trained without task tokens experience a notable increase in WER when a biasing list is used, indicating an inability to effectively utilize contextual information from the biasing list.
+Conversely, models trained with task tokens achieve a marked reduction in WER under both conditions, demonstrating that these tokens are essential for leveraging contextual cues and focusing on relevant entities from the biasing list.
+In Table II, we present representative results obtained from the two models discussed above.
+The absence of task tokens leads the model to overfit to the prompt, resulting in repeated or irrelevant words (e.g., coffee, no ) and reduced transcription accuracy.
+In contrast, incorporating task tokens enables the model to correctly identify and utilize relevant entities from the prompt (e.g., milk) while preserving accurate recognition of the remaining audio content.
 This example highlights that merely supplying prompt information is insufficient; without a multi-task framework and explicit task tokens, the model is prone to hallucinations and transcription errors.
@@
 0: Filtered list of entities
-1:  $hi$: index of the <hit> token
-2:  $H\leftarrow\text{Encoder}(\text{audio})$Compute encoder embedding}
+1:hihi: index of the <hit> token
+[…]
 3:for each entity $E$ in the biasing list do
@@
 9:return  Filtered list of entities
-Figure 2: Decoding process for the Prompt Biasing model.
-The model is trained to predict <hit> or <miss> for each sub-word in the prompt, enabling effective filtering of irrelevant entities.
-During decoding, entity filtering (see Algorithm 1) greatly reduces the biasing list size.
 III-B Entity Filtering for Biasing
 Previous studies on CLAS have shown that attention-based neural biasing methods encounter difficulties when processing large biasing lists [13].
-Our own experiments, as reflected in rows C1–C3 of Table II, corroborate this finding: large biasing lists introduce substantial noise, which diminishes the model’s ability to accurately focus on relevant entities.
+Our own experiments, as reflected in rows C1–C3 of Table III, corroborate this finding: large biasing lists introduce substantial noise, which diminishes the model’s ability to accurately focus on relevant entities.
 Our analysis reveals a clear correlation between the precision of the biasing list and the effectiveness of contextual biasing—the more targeted and relevant the biasing list, the greater the improvement in biasing performance.
@@
 The detailed steps of this entity filtering approach are provided in Algorithm 1.
-Although the method requires evaluating each entity in the biasing list, it is highly amenable to parallelization and can be executed with minimal latency on modern hardware accelerators such as GPUs.
-TABLE II: Performance on in-house domain dataset with various biasing lists.
+Although the method necessitates evaluating each entity in the biasing list, the additional computational overhead remains minimal.
+This efficiency is achieved by utilizing a shared encoder, eliminating the need for repeated computations.
+Furthermore, only a single forward pass through the relatively small decoder (approximately 100 million parameters) is required for each biasing entity, which can be efficiently batch-processed on modern hardware accelerators with minimal impact on latency.
+TABLE III: Performance on in-house domain dataset with various biasing lists.
 Row Model Biasing List EWER(%) WER(%)
@@
 To manage memory consumption, each training sample contains fewer than 20 entities in the biasing list.
-The model is trained on approximately 5,400 hours of anonymized paired audio-text data using the AdamW optimizer [34] and a linear decay learning rate schedule, with a peak learning rate of $2.24\times 10^{-4}$ and a brief warmup phase.
+The model was trained on roughly 5,400 hours of anonymized paired audio-text data sourced from various domains, such as voice assistants, conversational speech, and dictation tasks, ensuring a diverse dataset that reflects real-world usage.
+Model parameters are optimized using the AdamW optimizer [34] with a linear decay learning rate scheduler that features a peak learning rate of $2.24\times 10^{-4}$ and includes a short warmup period.
+TABLE IV: Number of unique entities in each domain dataset
+Domain Banking Healthcare Enhanced Medicine
+#Entity 973 1625 3143
+Domain Gaming Insurance Law Enforcement
+#Entity 1948 1296 1148
+Domain Basic Medicine Vehicles Science Technology
+#Entity 2633 1910 1584
 IV-B Evaluation Settings
@@
 Each utterance was annotated by human experts to identify entities of interest.
+Table IV presents the number of unique entities identified in each domain-specific datasets.
 To measure recognition performance on these entities, we utilized the Entity Word Error Rate (EWER), which specifically quantifies errors within labeled entities.
@@
 IV-C1 Contextual Biasing Performance
-Rows C1–C3 in Table II present the performance of the Prompt Biasing model evaluated with three different types of biasing lists on our in-house domain dataset.
+Rows C1–C3 in Table III present the performance of the Prompt Biasing model evaluated with three different types of biasing lists on our in-house domain dataset.
 First of all, when an ideal ground truth list (i.e.
 Exact list) is provided, the Prompt Biasing model achieves an impressive EWER of 1.8%, significantly outperforming its baseline model (i.e.
-A1 in Table II) by 76.8% relatively.
+A1 in Table III) by 76.8% relatively.
 This demonstrates the effectiveness of our approach in leveraging contextual information to enhance the entity recognition accuracy.
@@
 The Small list contains approximately 50 entities, while the Large list includes around 1800 entities.
-As shown in Table II (rows C2 and C3), the Prompt Biasing model achieves strong EWERs of 4.16% and 5.61% with the Small and Large lists, respectively, along with competitive WERs of 4.54% and 4.80%.
+As shown in Table III (rows C2 and C3), the Prompt Biasing model achieves strong EWERs of 4.16% and 5.61% with the Small and Large lists, respectively, along with competitive WERs of 4.54% and 4.80%.
 These results demonstrate that the model maintains effective contextual biasing and robust general transcription accuracy even when provided with imperfect biasing lists.
@@
 Given the observed benefits of providing a precise biasing list, we further evaluate the impact of our entity filtering strategy detailed in Algorithm 1.
-As shown in rows D1 and D2 of Table II, applying entity filtering leads to a notable improvement in EWER for both the Small and Large noisy biasing lists, compared to the standard prompt biasing model (rows C2 and C3).
+As shown in rows D1 and D2 of Table III, applying entity filtering leads to a notable improvement in EWER for both the Small and Large noisy biasing lists, compared to the standard prompt biasing model (rows C2 and C3).
 Specifically, EWER decreases from 4.16% to 4.04% for the Small list and from 5.61% to 4.95% for the Large list.
@@
 Shallow fusion is a widely adopted technique for enhancing contextual biasing performance in ASR systems.
-As shown in Table II, we compare our proposed Prompt Biasing model with entity filtering (rows D1 and D2) against the baseline model with shallow fusion (rows B1 and B2) across different biasing list sizes.
+As shown in Table III, we compare our proposed Prompt Biasing model with entity filtering (rows D1 and D2) against the baseline model with shallow fusion (rows B1 and B2) across different biasing list sizes.
 Our approach consistently outperforms shallow fusion, achieving a relative reduction in EWER of 30.7% for the Small biasing list (D1 vs.
@@
 IV-C4 Robustness Against Noise
-Table III presents the performance of the models on a 7.6M-word in-house dataset under both standard (no biasing list) and noisy biasing list conditions.
+Table V presents the performance of the models on a 7.6M-word in-house dataset under both standard (no biasing list) and noisy biasing list conditions.
 The standard setting corresponds to conventional ASR without contextual information.
@@
 Furthermore, when evaluated with a noisy biasing list, the Prompt Biasing model shows a negligible WER degradation of 0.06%, indicating strong robustness to irrelevant or inaccurate contextual information.
-TABLE III: Performance on in-house dataset for robustness benchmark.
+TABLE V: Performance on in-house dataset for robustness benchmark.
 Model Biasing List WER(%)